### v1

@@ Abstract
The Segment Anything Model (SAM) has gained significant attention for its impressive performance in image segmentation. However, it lacks proficiency in referring video object segmentation (RVOS) due to the need for precise user-interactive prompts and limited understanding of different modalities, such as language and vision. This paper presents the RefSAM model, which for the first time explores the potential of SAM for RVOS by incorporating multi-view information from diverse modalities and successive frames at different timestamps. Our proposed approach adapts the original SAM model to enhance cross-modality learning by employing a lightweight Cross-Modal MLP that projects the text embedding of the referring expression into sparse and dense embeddings, serving as user-interactive prompts. Subsequently, a parameter-efficient tuning strategy is employed to effectively align and fuse the language and vision features. Through comprehensive ablation studies, we demonstrate the practical and effective design choices of our strategy. Extensive experiments conducted on Ref-Youtu-VOS and Ref-DAVIS17 datasets validate the superiority and effectiveness of our RefSAM model over existing methods. The code and models will be made publicly at github.com/LancasterLi/RefSAM.
keywords:
[…]
1 Introduction
@@
Referring video object segmentation (RVOS) aims to segment the target object in the video with the guidance of given language expressions [1, 2]. It has attracted widespread attentions in the computer vision community due to its great potential in real world applications, such as video retrieval and video editing [3, 4, 5]. It is a challenging task and the key lies in how to incorporate multi-view information from video frames at different timestamps and align sources from different modalities, i.e., vision and language views. Compared with video object segmentation [6, 7], RVOS is more challenging as it is required to perform semantic-level alignment between vision and language modalities without the ground-truth mask annotation in the first frame. Although many efforts have been devoted to tackling this task [8, 9, 10, 11], it still remains a challenge how to accurately identify and segment the target object in the video by fully exploiting the semantic meanings of sources from vision and language views.
@@
Recently, the Segment Anything Model (SAM) [12] has been proposed, which serves as a foundation model for image segmentation. Based on the designed prompt engineering, it can be transferred to the new task in a zero-shot manner. Due to its impressive performance on image segmentation, many works have explored to apply SAM in other related fields [13, 14], e.g., medical image analysis [15, 16, 17], remote sensing images [18, 19, 20], video object tracking [21, 22], style transfer [23], and 3D reconstruction [24]. Although these methods have achieved great progress on these related tasks based on SAM, none of them can be applied to RVOS directly due to the inherent complexity of this task. Concretely, RVOS requires the model to have the comprehensive understanding of sources from vision and language modalities. Unlike video object segmentation (VOS), no ground-truth mask annotations in the first frame are available in RVOS. It’s worth noting that although some SAM-based multiple object tracking models such as SAM-Track [21] and TAM [22] can also be adapted to the task of RVOS by providing bounding boxes of the first frame detected by extra object detection models, such as Grounding DINO [25], some non-negligible issues can exist in such frameworks. For example, they cannot work effectively in an end-to-end manner due to the two-stage pipeline. As a result, inaccurate bounding boxes in the first frame detected by the object detection model can result in the poor performance of the downstream segmentation task. On the other hand, the involvement of separate models can bring extra difficulties in the model training and deployment process.
@@
Thus, a natural question arises that how to effectively adapt SAM to the RVOS task in an end-to-end manner to fully unleash its potential capacity for segmentation. In this paper, we conduct the initial exploration and propose the RefSAM, which is the first end-to-end SAM-based framework for the task of RVOS. Based on the powerful foundation model SAM, RefSAM can perform accurate target object segmentation in the video with given language expressions. Specifically, to enhance the cross-modality learning capability of the original SAM, we propose a lightweight Cross-Modal MLP that projects the text embedding of the referring expression into sparse and dense embeddings, serving as user-interactive prompts like points and bounding boxes prompts. Subsequently, a parameter-efficient tuning strategy is employed to effectively align and fuse the language and vision features.
@@ 1. $\bullet$
    We conduct the pioneering study to explore SAM’s potential for RVOS through the integration of multi-view information from diverse modalities and successive frames at different timestamps.
@@ 2. $\bullet$
    We introduce RefSAM, a novel approach that utilizes lightweight modules and an efficient fine-tuning strategy to align and fuse language and vision features in an end-to-end learning manner, thereby effectively adapting SAM for RVOS.
@@ 3. $\bullet$
    Extensive experiments on Ref-Youtu-VOS and Ref-DAVIS17 datasets demonstrate the promising results of RefSAM and its superiority compared to existing methods.
2 Related Work
2.1 Referring Video Object Segmentation
RVOS uses language cues to segment an object in a video. Unlike the VOS task which provides the ground truth bounding box of the first frame, it exploits a different type of guidance, i.e., language expressions, to identify and segment the object referred by the given language expression in a video. Therefore, the task is more challenging due to the great domain gap between language and vision modalities.
[26] first proposes the RVOS task and extends the A2D [27] and J-HMDB [28] datasets with sentences describing the actors and actions appearing in the video content. [29] extends language grounding models to video data to ensure temporally coherent predictions and augment Davis datasets [6, 30] with language descriptions. URVOS [31] proposes a unified end-to-end deep neural network that performs both language-based object segmentation and mask propagation in a single model and constructs a large-scale RVOS dataset Refer-Youtube-VOS. RefVOS [32] leverages the last stage feature of the vision backbone to fuse with the textual feature. ClawCraneNet [33] leverages three kinds of object-level relations to progressively construct discriminative visual embedding. CITD [2] puts forward a two-stage, top-down RVOS solution. MRSA [34] presents a multi-level visual representation framework that consists of video, frame, and object granularity. MMVT [9] incorporates the motion information from optical flow maps with appearance and linguistic features for text-based video segmentation. LBDT [8] proposes a Language-Bridged Duplex Transfer module to conduct spatial-temporal interaction explicitly between two independent 2D ConvNets in the encoding phase for RVOS. PMINet [10] performs multimodal interaction in each stage of the visual backbone, and multimodal feature is incorporated back into the visual backbone to guide the progressive learning of visual features. [11] utilizes implicitly individual word features to integrate visual features. These models still do not achieve outstanding performance due to the significant semantic gap between visual information and language information.
@@ 2.1 Referring Video Object Segmentation
Vision Transformers (ViTs) [35] have attracted lots of attention recently and have been widely used in various domains [36, 37, 38, 39, 40, 41], including the field of RVOS. MTTR [1] models the RVOS task as a parallel sequence prediction problem and outputs all objects in the video prior to selecting the one referred by the text. ReferFormer [4] introduces a small set of object queries that are conditioned on the text expression to attend to the referred object only. VLT [42] utilizes word features for query generation and spatial dynamic fusion. Different from these models, we present a simple yet effective solution by adapting SAM to RVOS via some minimal designs.
2.2 Segment Anything Model
@@
Recently, SAM [12] build a foundation model for segmentation. The model designs a promptable segmentation task and it can transfer zero-shot to new image distributions and tasks. Following that, a large number of research studies based on SAM emerged. SAM has been widely applied in various fields [13, 14], e.g., medical image analysis [15, 16, 17], remote sensing images [18, 19, 20], video object tracking [21, 22, 43, 44], style transfer [23], and 3D reconstruction [24]. In the medical image analysis domain, SAMed [45] involves customizing the SAM model for medical image segmentation by applying the low-rank-based (LoRA) fine-tuning strategy to the SAM image encoder which can perform better for semantic segmentation tasks on medical images. [46] demonstrates that SAM can achieve high segmentation accuracy for brain tumor MRI datasets in a point-to-mask setting. [47] assesses the zero-shot segmentation performance of the SAM model on representative segmentation tasks in whole slide imaging. In the remote sensing images domain, [19] introduces a domain-specific decoder, which learns the problem-specific semantics with only five labeled images. [20] develops an efficient pipeline with SAM to generate a large-scale remote sensing segmentation dataset named SAMRS. In the style transfer domain, Any-to-Any Style Transfer [48] enables users to specify which style region to select and which content regions to apply during style transfer. In the 3D reconstruction domain, SA3D [49] can segment any object in a 3D scene with one-shot manual prompting in a single rendered view. These advancements have shown the versatility of the foundation segmentation model.
PerSAM [50] proposes a training-free Personalization approach and a one-shot fine-tuning variant that only trains 2 parameters within 10 seconds for improved performance. HQ-SAM [51] designs a learnable High-Quality output token which injected into SAM’s mask decoder to predict the high-quality mask. These models all focus on the improvement of SAM in the vision field. SAM-Track [21] and TAM [22] enable users to select multiple objects in videos for tracking [52]. However, none of these models are currently available for the RVOS task. In this paper, we propose an end-to-end RVOS model based on SAM for the first time.
3 RefSAM for Reffering Video Object Segmentation
@@
First, we present the preliminaries of SAM in Sec. 3.1. Then, in Sec. 3.2, we provide an overview of the proposed RefSAM model. More details of each component of RefSAM are presented in the remaining sections.
@@
Figure 1: The overall pipeline of RefSAM. It mainly consists of five key components: Visual Encoder of SAM [12], Text Encoder, Cross Modal MLP, Dense Attention, Mask Decoder of SAM. We construct Sparse Embeddings and Dense Embeddings of cross-modal to learn text-visual information and predict masks.
3.1 Preliminaries
@@
Task Definition: Given a video clip $\mathcal{L}$ =$=\{I_{t}\}_{t=1}^{T}$ with $T$ frames and a referring expression $E=\{e_{l}\}_{l=1}^{L}$ with $L$ words, the goal of RVOS is to produce $T$-frame binary segmentation masks $S=\{s_{t}\}_{t=1}^{T}$, $s_{t}\in\mathbb{R}^{H\times W}$ of the referred object.
@@
Architecture of SAM: SAM [12] mainly consists of three components: an image encoder, a prompt encoder, and a mask decoder. The image encoder is a ViT-based [35] backbone for image feature extraction and can be applied prior to prompting the model. The prompt encoder encodes two sets of prompts, i.e., sparse (points and boxes) and dense (masks) prompts, to get interactive positional information and provide them to the mask decoder. The two-layer transformer-based mask decoder leverages the extracted image embedding, prompt embedding, and the learnable output and prompt tokens for final mask prediction. SAM demonstrates strong zero-shot generalization in the segmentation task. However, SAM is not skilled in utilizing text for segmentation and its training is very expensive. For example, training the ViT-H-based SAM model on SA-1B [12] requires 256 GPUs with a large batch size of 256 images.
3.2 Overview of RefSAM’s Architecture
@@
We introduce the RefSAM model to efficiently adapt SAM to the RVOS task and boost the potential segmentation capacity of SAM. As shown in Figure 1, it mainly consists of five key components: Visual Encoder, Text Encoder, Cross-Modal MLP, Dense Attention, and Mask Decoder. Firstly, we use the Visual Encoder of SAM to extract frame features as visual embeddings. Meanwhile, we use the text-to-text model (T5) [53] as the Text Encoder to extract linguistic embeddings. Then, we construct cross-modal Sparse Embeddings and Dense Embeddings to learn text-visual information and predict masks through the Cross-Modal MLP module. Next, RefSAM utilizes the Dense Attetion module to fuse visual embeddings and sparse embeddings in order to obtain dense embeddings. Finally, the Mask Decoder of SAM leverages sparse embeddings, dense embeddings, and visual embeddings for final mask prediction. During training, RefSAM reuses the pre-trained weights of SAM and text encoder T5 while only fine-tuning three modules, i.e., Cross-Modal MLP, Dense Attention, and Mask Decoder, enabling parameter-efficient fine-tuning. During inference, we directly output the mask predictions by selecting the masks with the highest score as the final results.
3.3 Backbone
3.3.1 Visual Encoder
@@
We start by adopting the image encoder of SAM as the visual encoder $\mathcal{E}_{v}$ to extract the visual feature maps for each frame in a video clip, as shown in the light blue part in the top of Figure 1. The image encoder $\mathcal{E}_{v}$ is an MAE [54] pre-trained ViT backbone [35]. Specifically, for each frame $I_{t}$ in the video clip $\mathcal{L}$ =$=\{I_{t}\}_{t=1}^{T}$, the vision encoder $\mathcal{E}_{v}$ is adopted to extract the feature map for this frame, i.e.,
$f_{t}=\mathcal{E}_{v}(I_{t})\in\mathbb{R}^{C_{v}\times H_{0}\times W_{0}},$ (1)
@@
where $f_{t}$ is the corresponding feature map of frame $I_{t}$. By applying the vision encoder $\mathcal{E}_{v}$ on each frame independently, a set of feature maps $\mathcal{F}_{v}=\{f_{t}\}_{t=1}^{T}$ can be obtained for $T$ frames in the video clip. As SAM has strong zero-shot segmentation performance, we freeze the parameters of the image encoder $\mathcal{E}_{v}$ to retain its feature extraction capability in the training process.
3.3.2 Text Encoder

### v3

@@ Abstract
The Segment Anything Model (SAM) has gained significant attention for its impressive performance in image segmentation. However, it lacks proficiency in referring video object segmentation (RVOS) due to the need for precise user-interactive prompts and a limited understanding of different modalities, such as language and vision. This paper presents the RefSAM model, which explores the potential of SAM for RVOS by incorporating multi-view information from diverse modalities and successive frames at different timestamps in an online manner. Our proposed approach adapts the original SAM model to enhance cross-modality learning by employing a lightweight Cross-Modal MLP that projects the text embedding of the referring expression into sparse and dense embeddings, serving as user-interactive prompts. Additionally, we have introduced the hierarchical dense attention module to fuse hierarchical visual semantic information with sparse embeddings to obtain fine-grained dense embeddings, and an implicit tracking module to generate a tracking token and provide historical information for the mask decoder. Furthermore, we employ a parameter-efficient tuning strategy to align and fuse the language and vision features effectively. Through comprehensive ablation studies, we demonstrate our model’s practical and effective design choices. Extensive experiments conducted on Refer-Youtube-VOS, Ref-DAVIS17, and three referring image segmentation datasets validate the superiority and effectiveness of our RefSAM model over existing methods. The code and models will be made publicly at github.com/LancasterLi/RefSAM.
Index Terms:
Video Object Segmentation, Vision Transformer, Language and Vision, Segment Anything
I Introduction
@@
Referring video object segmentation (RVOS) aims to accurately segment the target object in the video with the guidance of given language expressions [1, 2]. It has attracted widespread attention in the computer vision community due to its great potential in real-world applications, such as video retrieval and video editing [3, 4, 5]. It is a challenging task and the key lies in how to incorporate multi-view information from video frames at different timestamps and align sources from different modalities, i.e., vision and language views [6, 7]. Compared with video object segmentation, RVOS is more challenging as it is required to perform semantic-level alignment between vision and language modalities [8, 9] without the true mask labeling in the first frame. Although many efforts have been devoted to tackling this task [10, 11, 12], it still remains a challenge how to accurately identify and segment the target object in the video by fully exploiting the semantic meanings of sources from vision and language views.
@@
Recently, the Segment Anything Model (SAM) [13] has been proposed, which serves as a foundation model for image segmentation. Based on the designed prompt engineering, it can be transferred to the new task in a zero-shot manner. Due to its impressive performance on image segmentation, many works have explored to apply SAM in other related fields [14], e.g., medical image analysis [15], remote sensing images [16], video object tracking [17, 18], style transfer [19], and 3D reconstruction [20]. Although these methods have achieved great progress on these related tasks based on SAM, none of them can be applied to RVOS directly due to the inherent complexity of this task. Concretely, RVOS requires the model to have a comprehensive understanding of sources from vision and language modalities. Unlike video object segmentation (VOS), no ground-truth mask annotations in the first frame are available in RVOS. It’s worth noting that although some SAM-based multiple object tracking models such as SAM-Track [17] and TAM [18] can also be adapted to the task of RVOS by providing bounding boxes of the first frame detected by extra object detection models, such as Grounding DINO [21], some non-negligible issues can exist in such frameworks. For example, they cannot work effectively in an end-to-end manner due to the multi-stage pipeline. As a result, inaccurate bounding boxes in the first frame detected by the object detection model can result in the poor performance of the downstream segmentation task. On the other hand, the involvement of separate models can bring extra difficulties in the model training and deployment process.
@@
Thus, a natural question arises that how to effectively adapt SAM to the RVOS task in an end-to-end manner to fully unleash its potential capacity for video segmentation and multimodal fusion. In this paper, we conduct the initial exploration and propose the RefSAM, which is the first end-to-end SAM-based framework for the task of RVOS, as shown in Figure 1. Based on the powerful foundation model SAM and the specifically designed modules, RefSAM can perform accurate target object segmentation in the video with given language expressions in an online manner. The overall pipeline of RefSAM as depicted in Figure 2. Specifically, to enhance the cross-modality learning capability of the original SAM, we propose a lightweight Cross-Modal MLP that projects the text embedding of the referring expression into sparse and dense embeddings, serving as user-interactive prompts like points and bounding boxes prompts. Additionally, we have introduced the hierarchical dense attention module to fuse hierarchical visual semantic information with sparse embeddings in order to obtain fine-grained dense embeddings, and an implicit tracking module to generate a tracking token and provide historical information for the mask decoder. Furthermore, we employ a parameter-efficient tuning strategy to effectively align and fuse the language and vision features.
Figure 1: RefSAM integrates multi-view information from diverse modalities and successive frames at different timestamps in an online manner. Similar to SAM, RefSAM generates three types of token outputs: IOU token output (green), main mask token output (light blue), and three-scale mask token outputs (dark blue). We feed the main mask token output into RefSAM’s tracking module to obtain the track token that can provide historical information for the mask decoder and assist RefSAM in predicting the mask for the next frame.
@@ 1. $\bullet$
    We conduct the pioneering study to explore SAM’s potential for RVOS through the integration of multi-view information from diverse modalities and successive frames at different timestamps. Specifically, we introduce RefSAM, a novel approach that utilizes lightweight modules and an efficient fine-tuning strategy to align and fuse language and vision features in an end-to-end learning manner.
@@ 2. $\bullet$
    We design the hierarchical dense attention module for leveraging diverse levels of visual features and textual features to effectively perform cross-modal segmentation of objects with different sizes. We also devise an implicit tracking module to generate a track token and provide historical information to the mask decoder, which effectively enhances the spatiotemporal consistency of segmentation results.
@@ 3. $\bullet$
    Extensive experiments on Refer-Youtube-VOS, Ref-DAVIS17, and three Referring image segmentation (RIS) datasets demonstrate the promising results of RefSAM and its superiority compared to existing methods.
II Related Work
II-A Referring Video Object Segmentation
[22] first proposes the RVOS task, and presents an end-to-end pipeline that uses dynamic convolution for mask generation. Additionally, [22] extends the A2D-sentence [23] and J-HMDB [24] datasets with text descriptions. [25] first uses the grounding referring expression method to generate object boxes, and then utilizes a video segmentation model to create masks. [25] extends the DAVIS16 and DAVIS17 [26] datasets using two types of annotated language descriptions: “first frame” and “full video”. Later, URVOS [27] presents a large-scale RVOS dataset Refer-Youtube-VOS and proposes a new method that promotes cross-modal interactions. Many methods contribute to leveraging different levels of visual features to improve performance. RefVOS [28] utilizes the feature from the last stage of the encoder for cross-modal fusion. ClawCraneNet [29] explores the value of object-level information. MRSA [30] simultaneously acquires the video-level, frame-level, and object-level information during the encoding stage. Some other methods incorporate grounding techniques to assist in generating segmentation masks, such as CITD [31] and [12]. The former uses a grounding model for the final segmentation generation, while the latter employs one grounding model to select the best proposal for post-processing. Other approaches explore the utility of motion information. For example, MMVT [11] uses flow maps to capture motion information, while LBDT [10] relies on frame difference. Although these models have achieved promising performance, more research efforts are still needed to further address the significant semantic gap between visual information and language information.
@@ II-A Referring Video Object Segmentation
Vision Transformers (ViTs) [32] have attracted lots of attention recently and have been widely used in various domains [33, 34, 35, 36, 37, 38, 39], including the field of RVOS. MTTR [1] takes both a video and a text query as input, and it generates prediction sequences for all objects based on this information. It uses cross-entropy to align the video and text during the prediction process. ReferFormer [4] uses a set of language expression-based object queries as inputs to the Transformer decoder. These queries specifically focus on the referred object, and their purpose is to generate dynamic kernels. VLT [40] utilizes the QGM to produce various query vectors. These query vectors play a crucial role in guiding the Transformer decoder throughout the decoding process, ultimately generating appropriate responses. Different from these models, we present a simple yet effective solution by adapting SAM to RVOS via some minimal designs.
II-B Segment Anything Model
@@
Recently, SAM [13] builds a foundation model for segmentation. The model designs a promptable segmentation task and it can transfer zero-shot to new image distributions and tasks. Following that, a large number of research studies based on SAM emerged. SAM has been widely applied in various fields [14], e.g., medical image analysis [15], remote sensing images [16], video object tracking [17, 18], style transfer [19], and 3D reconstruction [20]. In the medical image analysis domain, SAMed [41] adds LoRA layers while freezing the image encoder to tune SAM for adapting to medical images. [42] demonstrates that SAM can achieve remarkable zero-shot segmentation accuracy on brain tumor MRI datasets. In the remote sensing images domain, [43] adds a domain decoder in SAM to learn specific features for the remote sensing images, while [16] leverages SAM to generate a large-scale dataset: SAMRS. In the style transfer domain, Any-to-Any Style Transfer [19] uses SAM to extract the segmentation mask and uses the VGG encoder and decoder to transform the segmentation mask into another image. In the 3D reconstruction domain, SA3D [44] employs a pre-trained NeRF and SAM to obtain 2D images, and utilizes inverse rendering and cross-view self-prompting to acquire corresponding 3D masks. These advancements have shown the versatility of the foundation segmentation model.
PerSAM [45] introduces target-guided attention and target-semantic prompting to enhance SAM’s performance in one-shot scenarios. Additionally, the model adaptively concatenates masks from three different scales, further improving its segmentation accuracy. HQ-SAM [46] incorporates a learnable HQ-token in SAM’s mask decoder to enhance the refinement of SAM’s output mask. SAM-Track [17] and TAM [18] enable users to select multiple objects in videos for tracking [47, 48, 49]. However, none of these models are currently available for the RVOS task. In this paper, we propose an end-to-end RVOS model based on SAM for the first time.
III RefSAM for Reffering Video Object Segmentation
@@
In this section, we introduce the RefSAM model. First, we present the preliminaries of SAM in Sec. III-A. Then, in Sec. III-B, we provide an overview of the proposed RefSAM model. Detailed descriptions of each component of RefSAM are provided in the subsequent sub-sections.
@@
Figure 2: The overall pipeline of RefSAM. It mainly consists of five key components: 1) Backbone: Visual Encoder of SAM [13] with Adapter and Text Encoder; 2) Cross-Modal MLP; 3) Hierarchical Dense Attention; 4) Mask Decoder of SAM; and 5) Implicit Tracking Module. We construct cross-modal Sparse Embeddings and Dense Embeddings to learn text-visual information and predict masks. We use the implicit tracking module to generate a track token and provide historical information for the mask decoder.
III-A Preliminaries
@@
Task Definition: The RVOS task leverages both visual and text information. Given a video clip $\mathcal{L}$ =$=\{I_{t}\}_{t=1}^{T}$ with $T$ frames and a referring expression $E=\{e_{l}\}_{l=1}^{L}$ with $L$ words, the goal of RVOS is to produce $T$-frame binary segmentation masks $S=\{s_{t}\}_{t=1}^{T}$, $s_{t}\in\mathbb{R}^{H\times W}$ of the referred object.
@@
Architecture of SAM: SAM [13] mainly consists of three components: an image encoder, a prompt encoder, and a mask decoder. The image encoder is a ViT-based [32] backbone to extract image features. The prompt encoder is responsible for encoding two sets of prompts: sparse prompts (consisting of points and boxes) and dense prompts (comprising masks). These prompts contain interactive positional information, which is then provided to the mask decoder for further processing. The mask decoder consists of a two-layer transformer. It takes image embeddings, output tokens, and prompt tokens as inputs and generates three-scale masks along with corresponding IoU scores as outputs. SAM demonstrates strong zero-shot generalization in the segmentation task. However, SAM has limitations in effectively leveraging text for segmentation, and its training process is computationally expensive due to the large scale of the SA-1B dataset and the huge amount of parameters.
III-B Overview of RefSAM’s Architecture
@@
We introduce the RefSAM model to efficiently adapt SAM to the RVOS task and boost the potential segmentation capacity of SAM. As shown in Figure 2, it mainly consists of five key components: two types of modality encoders, a Cross-Modal MLP, a Hierarchical Dense Attention (HDA) module, the Mask Decoder, and an Implicit Tracking Module (ITM). Firstly, we use the Visual Encoder of SAM to extract frame features as visual embeddings. Meanwhile, we use the text-to-text model (T5) [50] as the Text Encoder to extract linguistic embeddings. Then, we construct cross-modal Sparse Embeddings to provide aligned linguistic information to the Mask Decoder through the Cross-Modal MLP. Next, RefSAM utilizes the HDA module to fuse hierarchical visual semantic information and sparse embeddings in order to obtain fine-grained dense embeddings. Finally, the Mask Decoder of SAM leverages sparse embeddings, dense embeddings, and visual embeddings for final mask prediction and generates mask token output. The ITM utilizes the mask token output to generate a track token, which encodes historical information for the mask decoder. During training, RefSAM reuses the pre-trained weights of SAM and text encoder T5 while only fine-tuning lightweight modules, i.e., Adapter, Cross-Modal MLP, HDA, ITM, and Mask Decoder, enabling parameter-efficient fine-tuning. During inference, we directly output the mask predictions by selecting the masks with the highest score as the final results.
III-C Backbone
III-C1 Visual Encoder
@@
We start by adopting the image encoder of SAM as the visual encoder $\mathcal{E}_{v}$ to extract the visual feature maps and intermediate feature maps for each frame in a video clip, as shown in the blue part in the top of Figure 2. The image encoder $\mathcal{E}_{v}$ is an MAE [51] pre-trained ViT backbone [32]. Specifically, for each frame $I_{t}$ in the video clip $\mathcal{L}$ =$=\{I_{t}\}_{t=1}^{T}$, the vision encoder $\mathcal{E}_{v}$ is adopted to extract the feature maps Set for this frame, i.e.,
$f_{t},f_{t}^{mid}=\mathcal{E}_{v}(I_{t}),$ (1)
@@
where $f_{t}\in\mathbb{R}^{C_{v}\times H_{0}\times W_{0}}$ is the corresponding visual feature map of frame $I_{t}$ from the last layer of the SAM encoder. And $f_{t}^{mid}=\{f_{t}^{i}\}_{i=1}^{3}\in\mathbb{R}^{C_{mid}\times H_{0}\times W_{0}}$ denotes the set of intermediate feature maps of frame $I_{t}$. By applying the vision encoder $\mathcal{E}_{v}$ on each frame independently, a set of visual feature maps $\mathcal{F}_{v}=\{f_{t}\}_{t=1}^{T}$ and a set of intermediate feature maps $\mathcal{F}_{mid}=\{f^{mid}_{t}\}_{t=1}^{T}$ can be obtained for $T$ frames in the video clip. As SAM has strong zero-shot segmentation performance, we freeze the parameters of the image encoder $\mathcal{E}_{v}$ to retain its feature extraction capability in the training process.
Adapter in image encoder: Since SAM’s image encoder is adapted for purely visual tasks and does not align with text features, directly applying the frozen image encoder for RVOS leads to suboptimal outcomes. To improve the adaptability of the image encoder, we employ an adapter tuning approach [52, 53]. For the latter half of the transformer blocks in the SAM’s image encoder, each block is augmented with two trainable adapters. Each adapter consists of a linear layer for dimensionality reduction, a ReLU activation, and a linear layer for dimensionality expansion. We froze the image encoder of SAM and only fine-tuned the Adapter.
III-C2 Text Encoder